OpenStudy (anonymous):

x/4= (4x-3)/9.... solve for x.

OpenStudy (anonymous):

cross multiply 9x = 4(4x-3) 9x = 16x -12 7x = 12 x = 12/7

OpenStudy (anonymous):

ok with that Courtney?

OpenStudy (anonymous):

yes thank you